OpenStudy (ttop0816):

am i correct?? √(-1sin^2x)/(tan^2x+1) 0 1<< would it be this cos^2θ <

OpenStudy (anonymous):

???

OpenStudy (***[isuru]***):

is it... \[\frac{\sqrt{1 - \sin^{2} x} }{ \tan^{2} x +1 }\] ???

OpenStudy (ttop0816):

@***[ISURU]*** yes!